# Analysis of Variational Sparse Autoencoders

### v1

@@ Abstract
Sparse Autoencoders (SAEs) have emerged as a promising approach for interpreting neural network representations by learning sparse, human-interpretable features from dense activations. We investigate whether incorporating variational methods into SAE architectures can improve feature organization and interpretability. We introduce the variational Sparse Autoencoder (vSAE), which replaces deterministic ReLU gating with stochastic sampling from learned Gaussian posteriors and incorporates KL divergence regularization toward a standard normal prior. Our hypothesis is that this probabilistic sampling creates dispersive pressure, causing features to organize more coherently in the latent space while avoiding overlap. We evaluate a Topk vSAE against a standard TopK SAE on Pythia-70M transformer residual steam activations using comprehensive benchmarks including SAE Bench, individual feature interpretability analysis, and global latent space visualization through t-SNE. The vSAE underperforms standard SAE across core evaluation metrics, though excels at feature independence and ablation metrics. The KL divergence term creates excessive regularization pressure that substantially reduces the fraction of living features, leading to observed performance degradation. While vSAE features demonstrate improved robustness, they exhibit many more dead features than baseline. Our findings suggest that naive application of variational methods to SAEs does not improve feature organization or interpretability.
@@ 1.1 Background
Autoencoders were first applied to neural networks in 1986 by Rumelhart, Hinton, Williams. It is an architecture that learns to compress data into a smaller representation, then reconstructs the original data from that compressed form [1]. The network has two parts: an encoder that reduces input data to a lower-dimensional latent space, and a decoder that reconstructs the input from this compressed representation. The compressed middle layer forces the network to learn meaningful patterns and remove noise. [2]
@@
Variational autoencoders (VAEs) are probabilistic generative models that learn to encode data into a continuous latent space and decode it back to generate new samples. Unlike traditional autoencoders whose encoder maps inputs to fixed points in the latent space, encoder of the VAE instead produces two vectors: one for means and one for standard deviations, defining probability distributions for each latent dimension.
@@
Empirically, it is observed that autoencoders exhibit a non-smooth latent structure, while VAEs tend to have a smooth latent structure Latent Space Characterization of Autoencoder Variants. Research demonstrates that the latent spaces of standard autoencoders form non-smooth manifolds, while that of VAE forms a smooth manifold Latent Space Characterization of Autoencoder Variants. [5]
VAEs introduce a probabilistic structure to the latent space by learning a distribution over the latent variables rather than representing them as fixed points. This pushes the latent space toward continuous and smooth representations, facilitating tasks like data generation and interpolation with increased
@@ 1.2 Contributions
  • Comprehensive benchmarks using SAE Bench compare vSAE TopK against standard TopK SAE across multiple sparsity levels (k=64, 128, 256, 512) using the Pythia70m LM. We evaluate reconstruction fidelity, dictionary utilization, feature absorption, spurious correlation removal, and targeted probe perturbation to assess competitive performance.
@@ 1.3.2 SAEs
early work Using sparse representations to aid in the interpretability of neural networks dates back several decades. Early research involved the application of finding overcomplete sparse dictionary representations by the use of autoencoders on visual data. [13] [14] [15].
The later use of SAEs on LMs were found to produce more interpretable word vectors were found to be successful [16] [17] [18] [19]. Anthropic popularized the notion of features existing in a superposition across the neurons of a neural network [6], which led to the use of sparse autoencoders designed for the use of feature extraction to produce a monosemantic dictionary to represent the activations of a neural network [8] [9] [7].
@@ 4 Results
We would like to compare the vSAE against its SAE counterpart. We first look at the benchmark results to see how directly competitive the vSAE is. Then we look at the local, specific features which stood out to us. We then look at the global latent analysis of the features using tSNE.
@@ 4.1 Benchmarks
The difference between the two networks tested is the application of the kernel trick and the introduction of the KL loss instead of the traditional l2 loss used in the SAE. As the kernel trick is only a way to allow backpropagation in the network, we surmise that the dead features must be an effect of the KL loss term introduced to the vSAE. The pressure for activations to fit the gaussian footprint causes a quadratic loss pressure onto each feature instead of a linear pressure in the l2 of the SAE. Features unselected by the TopK mechanism still experience the pressure, meaning that many features will be penalized to the point that they stop activating altogether.
@@
The performance on the SCR benchmark agrees with our hypothesis: The SAE has spurious correlations that are deeply ingrained in the network, leading to degradation and performance loss when these connections are ablated. Our vSAE, on the other hand, experiences a more independent and less spuriously correlated feature space, so it maintains performance (or even slightly improves) when these spurious connections are removed. The two differences between the SAE and vSAE are the kernel trick and the KL divergence loss term. The kernel trick allows backpropagation while sampling from a distribution, so it is unlikely to be the cause of this change. This leaves the KL divergence term as the cause for the uplift in performance in this benchmark, meaning that the variational method has a more organized latent space capable of eschewing spurious correlation.
@@
  • The SAE peaks around a threshold of 8, while the vSAE has a higher peak near 64. A higher threshold is associated with sparsity, meaning that the SAE peaks while there are relatively more features active with generally lower activation values, and the vSAE peaks with fewer features but with stronger activations of those features. This implies that the vSAE latents are more robust and interpretable, and performs better when fewer features are active.
  • This graph introduces unintended effects to pair with the intended effects. We wish to minimize the unintended effects and maximize the intended. The SAE outperforms the vSAE on intended effects, but loses to the vSAE on the unintended effects. This implies that the vSAE is less likely to alter any of its representations if other representations ae ablated, while the SAE will be more severly changed in ways both beneficial and detrimental.
  • The high selectivity ratio in the vSAE means that ablating a class will have a small impact on other classes, implying a disentangled representation of the LM activation: the latent space is more dispersed and less entangled than the represenations of the SAE, and implies that the features of the vSAE are more aligned to the classes than the SAE. As the threshold grows, only the stronger latents with high feature activations for the class are still present, showing strong alignment.
@@ 4.2 Local
Although only 1,227 features were present in the vSAE histogram compared the SAE’s 6,970, we only saw a minor drop in performance. We also saw the vSAE outperform the SAE on the SCR and TPP benchmarks despite having fewer than one sixth of the total available features. The KL loss term has caused both an improvement in the independence of features, but has also deeply limited the computational capacity of the vSAE. The persistence of interpretability of the vSAE features is surprising given the small pool of options available to it, and brings into question the compressibility of model activations.
@@ 4.3 Global
Figure 10: The tSNE visualization of the SAE features. We can see that many features are active (1153 total), and a clear cluster is apparent in the middle right of the feature clusters graph. The feature utilization rate graph shows low utilization clustering in the bottom right, and the middle shows a collection of high-activation features. THis is seem as well in the third graph. The feature cluster utilization is quite varied.
Figure 11: The tSNE visualization of the vSAE features. Less clustering is immediately visible relative to the SAE features, and in the feature utilization rate we can see a greater dispertion of features with higher activations. We can also see a more equal distribution in the feature cluster utilization.
@@ 5 Conclusion
The variational Sparse Autoencoder confirms our central hypothesis: probabilistic sampling applied to a TopK sparse autoencoder creates dispersive pressure that organizes features more coherently in latent space. The benchmarks demonstrate this— our vSAE consistently outperforms standard SAE on spurious correlation removal and targeted probe perturbation, showing features with greater independence and less entanglement. The global t-SNE analysis further validates the hypothesis, revealing more uniform spatial distribution and balanced utilization patterns compared to the tightly clustered standard SAE representations. However, the addition of the KL divergence into the loss term also caused the majority of features to die in training, leading to worse performance than our baseline TopK sparse autoencoder in CE loss, explained variance, and cosine similarity tests.
These contradictory results reveal a fundamental trade-off: the KL regularization that improves feature independence simultaneously destroys representational capacity. The variational framework succeeds at its intended goal of creating dispersive pressure but at an unacceptable cost to dictionary utilization.

### v2

@@ Abstract
Sparse Autoencoders (SAEs) have emerged as a promising approach for interpreting neural network representations by learning sparse, human-interpretable features from dense activations. We investigate whether incorporating variational methods into SAE architectures can improve feature organization and interpretability. We introduce the Variational Sparse Autoencoder (vSAE), which replaces deterministic ReLU gating with stochastic sampling from learned Gaussian posteriors and incorporates KL divergence regularization toward a standard normal prior. Our hypothesis is that this probabilistic sampling creates dispersive pressure, causing features to organize more coherently in the latent space while avoiding overlap. We evaluate a TopK vSAE against a standard TopK SAE on Pythia-70M transformer residual stream activations using comprehensive benchmarks including SAE Bench, individual feature interpretability analysis, and global latent space visualization through t-SNE. The vSAE underperforms standard SAE across core evaluation metrics, though excels at feature independence and ablation metrics. The KL divergence term creates excessive regularization pressure that substantially reduces the fraction of living features, leading to observed performance degradation. While vSAE features demonstrate improved robustness, they exhibit many more dead features than baseline. Our findings suggest that naive application of variational methods to SAEs does not improve feature organization or interpretability.
@@ 1.1 Background
Autoencoders were first applied to neural networks in 1986 by Rumelhart, Hinton, and Williams. It is an architecture that learns to compress data into a smaller representation, then reconstructs the original data from that compressed form [1]. The network has two parts: an encoder that reduces input data to a lower-dimensional latent space, and a decoder that reconstructs the input from this compressed representation. The compressed middle layer forces the network to learn meaningful patterns and remove noise. [2]
@@
Variational autoencoders (VAEs) are probabilistic generative models that learn to encode data into a continuous latent space and decode it back to generate new samples. Unlike traditional autoencoders whose encoder maps inputs to fixed points in the latent space, the encoder of the VAE instead produces two vectors: one for means and one for standard deviations, defining probability distributions for each latent dimension.
@@
Empirically, it is observed that autoencoders exhibit a non-smooth latent structure, while VAEs tend to have a smooth latent structure Latent Space Characterization of Autoencoder Variants. Research demonstrates that the latent spaces of standard autoencoders form non-smooth manifolds, while those of VAEs form a smooth manifold Latent Space Characterization of Autoencoder Variants. [5]
VAEs introduce a probabilistic structure to the latent space by learning a distribution over the latent variables rather than representing them as fixed points. This pushes the latent space toward continuous and smooth representations.
@@ 1.2 Contributions
  • Comprehensive benchmarks using SAE Bench compare vSAE TopK against standard TopK SAE across multiple sparsity levels (k=64, 128, 256, 512) using the Pythia-70M LM. We evaluate reconstruction fidelity, dictionary utilization, feature absorption, spurious correlation removal, and targeted probe perturbation to assess competitive performance.
@@ 1.3.2 SAEs
Using sparse representations to aid in the interpretability of neural networks dates back several decades. Early research involved the application of finding overcomplete sparse dictionary representations by the use of autoencoders on visual data. [13] [14] [15].
The later use of SAEs on LMs were found to produce more interpretable word vectors successfully [16] [17] [18]. Anthropic popularized the notion of features existing in a superposition across the neurons of a neural network [6], which led to the use of sparse autoencoders designed for the use of feature extraction to produce a monosemantic dictionary to represent the activations of a neural network [8] [9] [7].
@@ 4 Results
We would like to compare the vSAE against its SAE counterpart. We first look at the benchmark results to see how directly competitive the vSAE is. Then we look at the local, specific features which stood out to us. We then look at the global latent analysis of the features using t-SNE.
@@ 4.1 Benchmarks
The difference between the two networks tested is the application of the reparameterization trick and the introduction of the KL loss instead of the traditional L2 loss used in the SAE. As the reparameterization trick is only a way to allow backpropagation in the network, we surmise that the dead features must be an effect of the KL loss term introduced to the vSAE. The pressure for activations to fit the Gaussian footprint causes a quadratic loss pressure onto each feature instead of a linear pressure in the L2 of the SAE. Features unselected by the TopK mechanism still experience the pressure, meaning that many features will be penalized to the point that they stop activating altogether.
@@
The performance on the SCR benchmark agrees with our hypothesis: The SAE has spurious correlations that are deeply ingrained in the network, leading to degradation and performance loss when these connections are ablated. Our vSAE, on the other hand, experiences a more independent and less spuriously correlated feature space, so it maintains performance (or even slightly improves) when these spurious connections are removed. The two differences between the SAE and vSAE are the reparameterization trick and the KL divergence loss term. The reparameterization trick allows backpropagation while sampling from a distribution, so it is unlikely to be the cause of this change. This leaves the KL divergence term as the cause for the uplift in performance in this benchmark, meaning that the variational method has a more organized latent space capable of eschewing spurious correlations.
@@
  • The SAE peaks around a threshold of 8, while the vSAE has a higher peak near 64. A higher threshold is associated with sparsity, meaning that the SAE peaks while there are relatively more features active with generally lower activation values, and the vSAE peaks with fewer features but with stronger activations of those features. This implies that the vSAE latents are more robust, interpretable, and perform better when fewer features are active.
  • This graph introduces unintended effects to pair with the intended effects. We wish to minimize the unintended effects and maximize the intended. The SAE outperforms the vSAE on intended effects, but loses to the vSAE on the unintended effects. This implies that the vSAE is less likely to alter any of its representations if other representations are ablated, while the SAE will be more severely changed in ways both beneficial and detrimental.
  • The high selectivity ratio in the vSAE means that ablating a class will have a small impact on other classes, implying a disentangled representation of the LM activation: the latent space is more dispersed and less entangled than the representations of the SAE, and implies that the features of the vSAE are more aligned to the classes than the SAE. As the threshold grows, only the stronger latents with high feature activations for the class are still present, showing strong alignment.
@@ 4.2 Local
Although only 1,227 features were present in the vSAE histogram compared to the SAE’s 6,970, we only saw a minor drop in performance. We also saw the vSAE outperform the SAE on the SCR and TPP benchmarks despite having fewer than one sixth of the total available features. The KL loss term has caused both an improvement in the independence of features, but has also deeply limited the computational capacity of the vSAE. The persistence of interpretability of the vSAE features is surprising given the small pool of options available to it, and brings into question the compressibility of model activations.
@@ 4.3 Global
Figure 10: The t-SNE visualization of the SAE features. We can see that many features are active (1153 total), and a clear cluster is apparent in the middle right of the feature clusters graph. The feature utilization rate graph shows low utilization clustering in the bottom right, and the middle shows a collection of high-activation features. This is seem as well in the third graph. The feature cluster utilization is quite varied.
Figure 11: The t-SNE visualization of the vSAE features. Less clustering is immediately visible relative to the SAE features, and in the feature utilization rate we can see a greater dispersion of features with higher activations. We can also see a more equal distribution in the feature cluster utilization.
@@ 5 Conclusion
We hypothesized that incorporating variational methods into sparse autoencoder architectures would create dispersive pressure through probabilistic sampling, organizing features more coherently and improving interpretability. Our results reject this hypothesis.
The variational Sparse Autoencoder exhibits catastrophic feature death, showing 82 percent fewer living features than standard TopK SAE driven by the KL divergence pressure that pushes activations toward zero. This capacity reduction causes degraded performance across core metrics: worse cross-entropy loss, lower explained variance, and reduced cosine similarity compared to baseline. While the vSAE demonstrates improvements in spurious correlation removal and targeted probe perturbation, these independence gains likely reflect reduced dictionary size rather than principled organization. Local feature analysis found no interpretability improvement despite substantially fewer active features.
Even under optimistic assumptions where feature death could be mitigated through architectural modifications, variational SAEs offer no practical advantage over standard approaches. The added computational overhead, training complexity, and hyperparameter sensitivity are not justified by marginal improvements in feature independence. We conclude that variational methods are not a promising direction for sparse autoencoder development.